OpenStudy (anonymous):

Could someone please help me with this? I'm not that great with angles What is the measure of each interior angle in a regular 15-gon? A.108° B.167° C.156° D.120°

OpenStudy (anonymous):

Equation: (n-2)180/15 n= the number of sides There is 2,340 degrees in a 15 gon, so you divide that by 15 which equals your answer

OpenStudy (anonymous):

The formula is (n-2)180/15 but thew answer is C 156 degrees

OpenStudy (anonymous):

If you need help in steps just ask

OpenStudy (anonymous):

actually the formula is \[\frac{(n-2)\times 180}{n}\]

OpenStudy (anonymous):

not \[\frac{(n-2)\times 180}{15}\]

OpenStudy (anonymous):

2340 / 15 is 156. So then the answer would be C? Thank you for your guys' help by the way!

OpenStudy (anonymous):

but in your case \(n=15\) and your answer is \[\frac{13\times 180}{15}\]